# Reference Points, Risk-Taking Behavior, and Competitive Outcomes in Sequential Settings

### v2

@@ 1 Introduction
To examine how pressure shapes the balance between ambition and feasibility in decision-making, we simulate counterfactual competition designs in which such pressure is removed. The results reveal consistent directional patterns—more conservative second attempts and more variable third attempts—but also considerable heterogeneity in both behavior and outcomes. While a subset of lifters benefit from the absence of pressure, many experience lower success rates and reduced expected outcomes, particularly in the second attempt. These differences reflect the uneven role of external pressure: for some, it enables focus and composure; for others, it disrupts execution or leads to overly cautious choices.
@@ 2.2 Game setting
The objective of each lifter is to lift the maximum weight possible from a prone position on a flat bench to improve their own rank. In case of a tie (same weight lifted), the lighter lifter or the lifter who achieved the weight first (depending on the federation’s rules) is ranked higher. Competitors are divided into categories based on weight class, age class, equipment class, and gender to ensure fair competition.
@@
Each lifter has three attempts.666 Lifting Procedure for each lifter at each attempt is officially divided into the following. 1. Setup: The lifter lies on their back on the bench, with feet flat on the floor or on the bench’s footrests, ensuring contact with the bench throughout the lift. 2. Grip: The lifter grips the barbell, usually slightly wider than shoulder-width apart. 3. Unracking: With the help of spotters if needed, the lifter unracks the barbell and holds it with arms fully extended above their chest. 4. Lowering: The barbell is lowered in a controlled manner to touch the chest, ensuring it is stable and paused for a moment. 5. Press Command: The judge gives a “Press” command once the barbell is motionless on the chest. The lifter then presses the barbell back to the starting position with arms fully extended. 6. Rack Command: Upon completion, and with the barbell in a stable position, the judge gives the “Rack” command, signaling the lifter to return the barbell to the rack. Each attempt is judged based on control, stability, and completion of lifting. A lift is considered successful if the lifter adheres to all the rules and completes the lift as per the judge’s commands from the viewpoints of two of three referees. After completing an attempt, the lifter must declare the weight for their next attempt within a specific time frame (usually 1 minute after their attempt). Once a weight is declared, it generally cannot be decreased, only increased. The order of lifters is adjusted after each round of attempts based on the declared weights for the next attempt, again proceeding from lightest to heaviest. This ensures that the competition is efficient and fair.
@@ 2.3 Exogenous pressure
In the second attempt of lifting the attempt weight, each lifter knows the first and second attempt outcomes of his one lower rank rival. These are uncorrelated with the lifter’s unobserved characteristics, such as his body condition on the competition date, because the lower-ranked rival completes both attempts without observing the lifter’s second outcome. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. If his second attempt fails, he is turned around by the rival. Second, each lifter also knows the second attempt weight of his one higher rank rival. This declared weight is uncorrelated with the lifter’s unobserved characteristics, because at the time of declaration, the higher-ranked rival has not yet observed the lifter’s second outcome. If his second attempt is successful, he has more chances to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain.
@@ IV X X
Overall, these findings highlight the evolving role of pressure across attempts, where loss aversion plays a role, and the difficulty of overtaking higher-ranked rivals remains a significant challenge.
@@ 5.3 Heterogeneity
Tables 7 and 8 show how responses to pressure vary by gender, experience, and historical rivalry frequency. Male lifters respond more strongly to both lower- and higher-ranked rivals (0.104 and 0.455, respectively) than female lifters (0.068 and 0.257), particularly under upward pressure. More experienced lifters are more responsive to higher-ranked rivals (0.031–0.032), reinforcing the role of competitive cues accumulated through experience. A higher historical rivalry frequency, defined as the number of prior encounters with a rival, is associated with greater responsiveness (0.100 and 0.095 for second attempts; 0.127 and 0.095 for third attempts), suggesting that familiarity with rivals plays a role in shaping attempt selection. These trends indicate that male experienced lifters facing well-known rivals take greater risks.
For success probabilities, gender differences vary depending on the direction of pressure. When lifters are under pressure from lower-ranked rivals—i.e., at risk of being overtaken—male lifters consistently show higher success probabilities in both the second (0.015) and third attempts (0.032). Female lifters, by contrast, show little to no benefit in the second attempt (–0.003), but exhibit a positive response in the third attempt (0.034), suggesting that the effect of downward pressure converges across genders in later stages.
In contrast, when attempting to overtake higher-ranked rivals, both male and female lifters experience reduced success probabilities, with effects slightly more negative for female lifters (–0.039 vs. –0.034 in the third attempt). These results indicate that upward pressure imposes psychological and performance costs for most lifters, with gender gaps persisting, while the motivational effects of downward pressure appear more stage-dependent and gender-convergent.
Experience plays a limited role in moderating these effects, with small but statistically significant coefficients. Historical rivalry frequency is associated with higher success probabilities when under pressure from lower-ranked rivals (0.012 in the second attempt and 0.028 in the third), but with lower success probability when facing pressure from higher-ranked rivals (–0.007 in the third attempt). This pattern suggests that familiarity with lower-ranked rivals may help lifters manage performance anxiety, while familiarity with higher-ranked rivals may instead heighten psychological stress. These results emphasize the complex interplay between gender, experience, and competitive history in shaping both risk-taking behavior and execution under pressure.
@@ 6 Counterfactual
Using the estimated results, we first simulate the attempt weights relative to the current best weight and the success probability for each attempt of each player under actual pressure. We then calculate the expected achieved weights relative to the current best, defined as the product of the attempt weight minus the current best weight and the success probability. These serve as benchmark outcomes under actual pressure conditions.
@@
For these alternative competition settings, we simulate the attempt weights relative to the current best weight, the success probability, and the expected achieved weight for each player, holding the estimated coefficients of the pressure variables at zero.
@@ 6.1 Attempt and success probability
Panel (a) in Figure 4 shows the distribution of proportional changes in second and third attempt weights relative to the current best under the counterfactual without pressure during the attempt. In the second attempt, most lifters select lighter weights than in the benchmark, resulting in a left-skewed distribution. In the third attempt, the distribution is more symmetric and concentrated near zero, with modest variation in both directions. These patterns suggest that pressure has a stronger and more uniform effect on second attempts than third attempts.
@@
Panel (b) in Figure 4 presents the distribution of proportional changes in second and third attempt success probabilities from the benchmark to the counterfactual. For the second attempt, the distribution is relatively concentrated, with peaks around zero, indicating that some lifters experience modest gains or losses in success probability when pressure is removed. In contrast, the third attempt shows a more dispersed distribution with heavier tails, suggesting greater variability in how lifters respond to the absence of pressure. While a subset of lifters benefit, many experience reduced success. Overall, the removal of competitive pressure leads to mixed effects in both attempts, with more pronounced heterogeneity in the third.
@@ 6.2 Expected weight
Figure 5: Counterfactual Expected Achieved Weight Minus Current Best Weight
@@
Figure 5 presents the distribution of expected achieved weight differences from the lifters’ current best for second and third attempts, comparing the conditions of “no pressure” during the attempt, lifting, or both. We take the differences to incorporate individual-level heterogeneity on the best achieved weight.
In the second attempt, all counterfactual settings lead to lower expected outcomes, especially in the “no pressure in both” condition, where the distribution is clearly left-skewed. The “no pressure during lifting” case shows a sharper peak near zero, suggesting that many lifters would still choose similar weights without pressure during execution.
@@
As practical advice, most lifters may benefit from adjusting their second attempts in response to competitive pressure, as removing pressure tends to reduce both attempted weight and success probability. In contrast, maintaining one’s original plan in the third attempt—rather than reacting to pressure from rivals—may help preserve focus, though this strategy is not universally optimal across lifters. More broadly, individuals in sequential high-stakes settings should account for their own sensitivity to competitive pressure, recognizing that for many, performance deteriorates in its absence.
@@ IV X X
Table 9 corresponds to Table 5 for the analysis of attempt weights using the lifter’s current best within one year. The main results remain robust under this alternative specification. Male lifters continue to exhibit stronger responses in attempt weight selection, with coefficients of 0.528 statistically insignificant for the second attempt and 5.785 for the third attempt. Body weight remains positively associated with attempt weights (0.151 and 0.254), while experience shows a small negative effect (–0.059 and –0.125), suggesting that more experienced lifters may exhibit slightly more conservative choices.
The coefficients for pressure from lower- and higher-ranked rivals remain positive and statistically significant. For the second attempt, the effects are 0.092 (lower rival) and 0.483 (higher rival), while for the third attempt, the corresponding coefficients are 0.043 and 0.427. These patterns reinforce the idea that lifters are sensitive to competitive dynamics, particularly when higher-ranked rivals lift successfully, leading to upward adjustments in attempt weights.
@@ 8 Conclusion
Counterfactual simulations show that without competitive pressure, lifters tend to adopt more conservative second attempts and maintain or slightly increase third attempt weights. However, the effects on performance are highly heterogeneous: while some lifters benefit from the absence of pressure, many experience lower success rates and reduced expected outcomes, particularly in the second attempt. These results highlight that external competition plays a dual role—motivating some lifters to perform better while destabilizing others—underscoring the importance of individual sensitivity in shaping performance under pressure.

### v3

@@ 1 Introduction
To examine how pressure shapes the balance between ambition and feasibility in decision-making, we simulate counterfactual competition designs in which such pressure is removed. The results reveal consistent directional patterns—more conservative attempts—but also considerable heterogeneity in expected outcomes. While a subset of lifters benefit from the absence of pressure, some lifters experience lower success rates and reduced expected outcomes. These differences reflect the uneven role of external pressure: for some, it enables focus and composure; for others, it disrupts execution or leads to overly cautious choices.
@@ 2.2 Game setting
The objective of each lifter is to lift the maximum weight possible from a prone position on a flat bench to improve their own rank. In case of a tie (same weight lifted), the lighter lifter or the lifter who achieved the weight first (depending on the federation’s rules) is ranked higher. Competitors are divided into categories based on weight class, age class, equipment class, and gender to ensure fair competition. [^6] See details in the Internatinal Powerlifting Federation (IPF) Technical Rules Book.
@@
Each lifter has three attempts.777 Lifting Procedure for each lifter at each attempt is officially divided into the following. 1. Setup: The lifter lies on their back on the bench, with feet flat on the floor or on the bench’s footrests, ensuring contact with the bench throughout the lift. 2. Grip: The lifter grips the barbell, usually slightly wider than shoulder-width apart. 3. Unracking: With the help of spotters if needed, the lifter unracks the barbell and holds it with arms fully extended above their chest. 4. Lowering: The barbell is lowered in a controlled manner to touch the chest, ensuring it is stable and paused for a moment. 5. Press Command: The judge gives a “Press” command once the barbell is motionless on the chest. The lifter then presses the barbell back to the starting position with arms fully extended. 6. Rack Command: Upon completion, and with the barbell in a stable position, the judge gives the “Rack” command, signaling the lifter to return the barbell to the rack. Each attempt is judged based on control, stability, and completion of lifting. A lift is considered successful if the lifter adheres to all the rules and completes the lift as per the judge’s commands from the viewpoints of two of three referees. After completing an attempt, the lifter must declare the weight for their next attempt within a specific time frame (usually 1 minute after their attempt). Once a weight is declared, it generally cannot be decreased, only increased. The order of lifters is adjusted after each round of attempts based on the declared weights for the next attempt, again proceeding from lightest to heaviest. This ensures that the competition is efficient and fair.
@@ 2.3 Exogenous pressure
In the second attempt of lifting the attempt weight, each lifter knows the first and second attempt outcomes of his one lower rank rival. These are uncorrelated with the lifter’s unobserved characteristics, such as his body condition on the competition date, because the lower-ranked rival completes both attempts without observing the lifter’s second outcome. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. If his second attempt fails, he is turned around by the rival. Second, each lifter also knows the second attempt weight of his one higher rank rival. This declared weight is uncorrelated with the lifter’s unobserved characteristics, because at the time of declaration, the higher-ranked rival has not yet observed the lifter’s second outcome. If his second attempt is successful, he has more chances to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain. Note that the pressure matters only in the lifting stage.
@@ IV X X
Overall, these findings highlight the role of pressure across attempts, with loss aversion playing a key role and overtaking higher-ranked rivals remaining particularly challenging. The asymmetry—where upward pressure consistently lowers success probabilities, while downward pressure has neutral or slightly positive effects—is consistent with the loss-averse utility structure assumed in the theoretical framework, whereby individuals exert more effort under pressure from lower-ranked rivals to avoid losses. [^13] This pattern is also observed in other competitive settings, such as marathons (Allen et al. 2017).
@@ 5.3 Heterogeneity
Tables 7 and 8 show how responses to pressure vary by gender, experience, and historical rivalry frequency. Male lifters respond more strongly to both lower- and higher-ranked rivals (0.104 and 0.455, respectively) than female lifters (0.068 and 0.257), particularly under upward pressure. More experienced lifters are more responsive to higher-ranked rivals, highlighting the role of accumulated competitive cues. A higher historical rivalry frequency—defined as the number of prior encounters with a rival—is associated with greater responsiveness, suggesting that familiarity influences attempt selection. These trends indicate that male, experienced lifters facing well-known rivals take greater risks.
For success probabilities, gender differences vary by pressure direction. When under pressure from lower-ranked rivals—i.e., at risk of being overtaken—male lifters show higher success probabilities in both second (0.015) and third attempts (0.032). Female lifters, by contrast, show little to no benefit in the second attempt (–0.003) but exhibit a positive effect in the third (0.034), suggesting that downward pressure converges across genders in later stages.
In contrast, when attempting to overtake higher-ranked rivals, both genders experience reduced success probabilities, slightly more negative for female lifters (–0.039 vs. –0.034 in the third attempt). This suggests that upward pressure imposes psychological and performance costs for most lifters, with gender gaps persisting, while motivational effects of downward pressure are more stage-dependent and gender-convergent.
Experience plays a limited role in moderating these effects, with small but significant coefficients. Historical rivalry frequency is linked to higher success under downward pressure (0.012 in the second attempt; 0.028 in the third), but lower success under upward pressure (–0.007 in the third). This suggests familiarity with lower-ranked rivals helps manage performance anxiety, whereas familiarity with higher-ranked rivals may heighten psychological stress. Overall, these results underscore the complex interplay between gender, experience, and competitive history in shaping risk-taking and execution under pressure.
@@ 6 Counterfactual
Using the estimated results, we first simulate the attempt weights and the success probability for each attempt of each player under actual pressure. [^14] Our regression results use the predicted attempt weight minus the current best as a dependent variable. We obtain simulated attempt weight by adding the current best to the predicted outcomes. We then calculate the expected achieved weights, defined as the product of the predicted attempt weight and the success probability. These serve as benchmark outcomes under actual pressure conditions.
@@
For these alternative competition settings, we simulate the attempt weights, the success probability, and the expected achieved weight for each player, holding the estimated coefficients of the pressure variables at zero.
@@ 6.1 Attempt and success probability
Panel (a) in Figure 4 shows the distribution of proportional changes in second and third attempt weights under the counterfactual without pressure during the attempt. Both distributions exhibit substantial deviations from zero, indicating that pressure has a strong influence on attempt weights. The third attempt distribution is slightly more concentrated around zero, but the overall shapes of the second and third attempt distributions are similar. These patterns suggest that pressure during the attempt significantly affects weight selection, with comparable impacts on second and third attempts.
@@
Panel (b) in Figure 4 presents the distribution of proportional changes in the second and third attempt success probabilities from the benchmark to the counterfactual. For the second attempt, the distribution is relatively concentrated, with peaks around zero, indicating that some lifters experience modest gains or losses in success probability when pressure is removed. In contrast, the third attempt shows a more dispersed distribution with heavier tails, suggesting greater variability in how lifters respond to the absence of pressure. While a subset of lifters benefit, many experience reduced success. Overall, the removal of competitive pressure leads to mixed effects in both attempts, with more pronounced heterogeneity in the third.
@@ 6.2 Expected weight
Figure 5: Counterfactual Expected Achieved Weight
@@
Figure 5 presents the distribution of expected achieved weight differences for second and third attempts, comparing the conditions of “no pressure” during the attempt, lifting, or both.
In the second attempt, almost all counterfactual settings lead to lower expected outcomes, especially in the “no pressure in both” condition, where the distribution is clearly left-skewed. The “no pressure during lifting” case shows a sharper peak near zero, suggesting that many lifters would still achieve similar weights without pressure during execution.
@@
As practical advice, most lifters may benefit from adjusting their attempts in response to competitive pressure, as removing pressure tends to reduce both attempted and achieved expected weights. However, reacting to pressure from rivals is not universally optimal across lifters. More broadly, individuals in sequential high-stakes settings should account for their own sensitivity to competitive pressure, recognizing that for many, performance deteriorates in its absence.
@@ IV X X
Table 9 corresponds to Table 5 for the analysis of attempt weights using the lifter’s current best within one year. The main findings remain robust: pressure from both lower- and higher-ranked rivals continues to have positive and significant effects on attempt weights in both the second and third attempts. While the coefficient on the male dummy changes sign in the second attempt, the overall patterns are consistent, reinforcing that lifters respond to competitive pressure, particularly from higher-ranked rivals.
@@ 8 Conclusion
Counterfactual simulations show that without competitive pressure, lifters tend to adopt more conservative attempts. However, the effects on performance are highly heterogeneous: while some lifters benefit from the absence of pressure, some lifters experience lower success rates and reduced expected outcomes. These results highlight that external competition plays a dual role—motivating some lifters to perform better while destabilizing others—underscoring the importance of individual sensitivity in shaping performance under pressure.